if n is divisible by 24 and 36 than how it would be divisible by 18
Answer by algebrahouse.com(1659)

"if n is divisible by 24 and 36 than how it would be divisible by 18"

If n = 72
- it is divisible by all three numbers

72 ÷ 24 = 3
72 ÷ 36 = 2
72 ÷ 18 = 4
www.algebrahouse.com